### v1

# Smallness of neutrino masses and leptogenesis in

Abstract
We consider 331 composite Higgs model (CHM3) in which the Lagrangian of the strongly coupled sector is invariant with respect to global $SU(3)_{C}\times SU(3)\times U(1)_{6}$ symmetry that can originate from $SU(6)$ subgroup of $E_{6}$ and contains the gauge group of the standard model (SM) as a subgroup. The breakdown of the approximate $SU(3)\times U(1)_{6}$ symmetry down to $SU(2)_{W}\times U(1)_{Y}$ subgroup around the scale $f\sim 10\,\mbox{TeV}$ results in a set of pseudo–Nambu–Goldstone bosons (pNGBs) that, in particular, involves Higgs doublet. The generation of the masses of the SM fermions in the CHM3 is discussed. We argue that an approximate discrete $Z_{2}$ symmetry may give rise to tiny masses of the left–handed neutrinos and several composite fermions with masses $1-2\,\mbox{TeV}$. The lepton and baryon asymmetries can be generated within the CHM3 via the out–of equilibrium decays of extra Majorana particle into the Higgs doublet and these composite fermions.
1 Introduction
The smallness of three neutrino masses may be naturally explained within seesaw models [1]. In these models the lepton asymmetry can be induced via the out–of equilibrium decays of the lightest right–handed neutrino [2] since all three Sakharov conditions [3] are satisfied in this case. This asymmetry gets partially converted into baryon asymmetry via sphaleron processes [4].
In this context it is especially interesting to consider the possible origin of the mass hierarchy in the lepton sector and leptogenesis within well motivated extensions of the standard model (SM) that allow to almost stabilize the electroweak (EW) scale such as composite Higgs models (CHMs). These extensions of the SM proposed in the 70’s [5] and 80’s [6] involve two sectors. Apart from the weakly–coupled sector, that includes elementary states with the quantum numbers of the SM fermions and SM gauge bosons, there are their composite partners which are formed in the second strongly interacting sector [7]. The Higgs doublet in these models is a set of bound states.
The Lagrangian of the strongly interacting sector of the minimal composite Higgs model (MCHM) [8, 9] possesses an approximate global $SU(3)_{C}\times SO(5)\times U(1)_{X}$ symmetry. $SU(3)_{C}\times SU(2)_{W}\times U(1)_{Y}$ gauge group of the SM is a subgroup of $SU(3)_{C}\times SO(5)\times U(1)_{X}$. Near the scale $f$ the approximate $SO(5)\times U(1)_{X}$ symmetry is expected to be broken down to $SO(4)\times U(1)^{\prime}_{X}\cong SU(2)_{W}\times SU(2)_{R}\times U(1)^{\prime}_{X}$. Such breakdown results in four pseudo–Nambu–Goldstone bosons (pNGBs) that compose the Higgs doublet $H$. The custodial symmetry $SU(2)_{cust}\subset SO(4)\cong SU(2)_{W}\times SU(2)_{R}$ [10] protects the Peskin–Takeuchi $\hat{T}$ parameter [11] against new physics contributions. The contributions of new bound states to the EW observables in the CHMs were analysed in Refs. [12][20]. The implications of these models were also examined for Higgs physics [15][16], [21][24], gauge coupling unification [25][26], dark matter [13], [22], [26][27] and collider phenomenology [14][15], [17], [24], [28][30]. Different aspects of non-minimal CHMs were explored in Refs. [13], [21][22], [26][27], [31].
The SM bosons and fermions except the Higgs doublet are superpositions of elementary states and their composite partners. Within partial compositeness framework [32, 33] the couplings of the SM states to the Higgs doublet are determined by the fractions of their compositeness. The mass hierarchy in the quark and charged lepton sectors can be reproduced in this case if the fractions of compositeness of all fermions except the top quark are rather small. This also leads to the partial suppression of flavour–changing processes [32]. The non–diagonal flavour transitions in the quark and lepton sectors within the CHMs were examined in Refs. [18][20], [28], [34][35] and [29], [35][37], respectively. It was shown that in general the compositeness scale $f$ in these models is required to be larger than $10\,\mbox{TeV}$ [18][19], [28], [34], [36]. However in the CHMs with extra flavour symmetries the phenomenologically acceptable scenarios can be obtained even for $f\sim 1\,\mbox{TeV}$ [17][18], [28][29], [35], [38].
In this note the mass hierarchy in the lepton sector and leptogenesis are discussed in the framework of 331 composite Higgs model (CHM3). The strongly interacting sector of the CHM3 possesses a global $SU(3)_{C}\times SU(3)\times U(1)_{6}$ symmetry which contains the SM gauge group as a subgroup and may stem from $SU(6)$ subgroup of $E_{6}$. The $E_{6}$ inspired composite Higgs models (E6CHM) were explored in Refs. [39, 40] [^2] The $E_{6}$ inspired supersymmetric extensions of the SM were studied in [41].. Near the compositeness scale $f$ the breakdown of the approximate $SU(3)\times U(1)_{6}$ symmetry takes place giving rise to the composite Higgs doublet. Since the $SU(3)$ symmetry does not contain $SU(2)_{cust}$ subgroup the electroweak precision measurements imply that $f\gtrsim 5-6\,\mbox{TeV}$ [39].
The paper is organised as follows. In the next section we specify the CHM3, discuss the generation of masses of the SM fermions and argue that an approximate discrete $Z_{2}$ symmetry can lead to tiny masses of the left–handed neutrino. This discrete symmetry may also result in relatively light composite partners of the left–handed leptons which might be detected at the LHC in the near future. The leptogenesis in such scenario is considered in section 3. Our results are summarised in section 4.
2 CHM3 with approximate $Z_{2}$ symmetry
The measured Higgs mass $m_{h}\approx 125\,\mbox{GeV}$ corresponds to a quite small value of the Higgs quartic coupling, $\lambda\simeq 0.13$ in the SM Higgs potential. On the other hand in general $\lambda$ tends to be of the order of unity in the CHMs. The relatively small values of $m_{h}$ and $\lambda$ indicate that the Higgs doublet can appear as a set of pNGB states from the breaking of an approximate global symmetry of the second strongly coupled sector. The couplings of the elementary states to the fields of the strongly interacting sector explicitly break this global symmetry. The pNGB potential arises from loops that contain elementary states. Therefore the quartic coupling $\lambda$ in the Higgs effective potential is somewhat suppressed.
Within the CHMs it is convenient to present the pNGB states as
$\Sigma=e^{i\Pi/f}\,,\qquad\Pi=\Pi^{\hat{a}}T^{\hat{a}}\,$, (1)
where $T^{\hat{a}}$ are broken generators of the global symmetry. The breaking $SU(3)\times U(1)_{6}\to SU(2)_{W}\times U(1)_{Y}$ can be parameterised through the fundamental representation of $SU(3)$, i.e.
$\Omega=\Sigma\,\Omega_{0}\,,\qquad\qquad\Omega_{0}^{T}=(0\quad 0\quad 1)\,.$ (2)
When $f$ is much larger than the vacuum expectation value of the Higgs field $v\simeq 246\,\mbox{GeV}$ the first two components of $\Omega$ can be identified with the SM–like Higgs doublet while the third component involves the SM singlet scalar $\phi_{0}$. In the CHM3 the $U(1)_{Y}$ weak hypercharges $Y_{i}$ are linear combinations
$Y_{i}=\frac{2\sqrt{3}}{6}T^{8}_{i}+Q^{6}_{i}\,$, (3)
where $Q^{6}_{i}$ are $U(1)_{6}$ charges of different multiplets and in the $SU(3)$ fundamental representation
$T^{8}=\dfrac{1}{2\sqrt{3}}\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&0&-2\end{array}\right)\,.$
To ensure the appropriate breakdown of $SU(3)\times U(1)_{6}$ symmetry $\Omega$ has to carry $U(1)_{6}$ charge (+1/3).
In the MCHM all composite partners of the SM fermions can be embedded into the vector representations of $SO(5)$ [9], i.e. $\bf 5$}–plets. This $SO(5)$ representation has the following decomposition in terms of $SU(2)_{W}\times SU(2)_{R}$:
${\bf 5}=({\bf 2,2})\oplus{\bf 1}\,.$ (4)
In Eq. (4) the first and second quantities in brackets are the $SU(2)_{W}$ and $SU(2)_{R}$ representations. In particular, the composite partners of the left–handed and charged right–handed leptons ($L^{\prime}_{1i}$ and $E^{{}^{\prime}c}_{i}$) may belong to ${\bf 5}^{i}_{-1}$ and ${\bf 5}^{i}_{+1}$ respectively where lower index corresponds to the $U(1)_{X}$ charge of the 5–plet whereas $i=1,2,3$ runs over all three generations. However the interactions $(L^{\prime}_{1i}H)(L^{\prime}_{1j}H)$, that could generate the masses of the left–handed neutrino, are forbidden in this model by the $U(1)_{X}$ symmetry. Such interactions may appear if another composite partners $L^{\prime}_{2i}$ of the left–handed leptons, which are components of ${\bf 5}^{i}_{0}$, are included. The ${\bf 5}^{i}_{0}$ contain the following $SU(2)_{W}\times U(1)_{Y}$ multiplets
$X_{i}=\left({\bf 2},\,+\dfrac{1}{2}\right)\,,\qquad\qquad L^{\prime}_{2i}=\left({\bf 2},\,-\dfrac{1}{2}\right)\,,\qquad\qquad N_{i}=({\bf 1},\,0)\,.$ (5)
where the first and second numbers in brackets are the $SU(2)_{W}$ representation and $U(1)_{Y}$ charge. The composite partners of the right–handed up– and down–quarks ($U^{{}^{\prime}c}_{i}$ and $D^{{}^{\prime}c}_{i}$) are components of ${\bf 5}^{i}_{-2/3}$ and ${\bf 5}^{i}_{+1/3}$ respectively. There are two types of the partners of the left–handed quarks ($Q^{\prime}_{1i}$ and $Q^{\prime}_{2i}$) in this case that belong to ${\bf 5}^{i}_{+2/3}$ and ${\bf 5}^{i}_{-1/3}$.
In the simplest $SU(5)$ Grand Unified Theories (GUTs) the left-handed quark and lepton doublets ($q_{i}$ and $\ell_{i}$), the right-handed charged leptons ($e_{i}^{c}$), the right-handed up- and down-type quarks ($u_{i}^{c}$ and $d_{i}^{c}$) are components of the $SU(5)$ multiplets, i.e.
$u^{c}_{i}\in{\bf 10}_{i}\,,\qquad q_{i}\in{\bf 10}_{i}\,,\qquad d^{c}_{i}\in{\bf\overline{5}}_{i}\,,\qquad e^{c}_{i}\in{\bf 10}_{i}\,,\qquad\ell_{i}\in{\bf\overline{5}}_{i}\,.$ (6)
The Lagrangian of the strongly interacting sector of the CHM3 is invariant under the transformations of the global $SU(3)_{C}\times SU(3)\times U(1)_{6}$ symmetry which is a subgroup of $SU(6)$. The normalisation of the $U(1)_{6}$ charges used in this paper implies that the fundamental representation of the $SU(6)$ group involves
${\bf 6}=\left({\bf 3},\,{\bf 1},\,-\dfrac{1}{3}\right)\oplus\left({\bf 1},\,{\bf 3},\,\dfrac{1}{3}\right)\,.$ (7)
Hereafter the first, second and third quantities in brackets are the $SU(3)_{C}$ and $SU(3)$ representations as well as $U(1)_{6}$ charges. The $\bf 5$}–plet and $\bf 10$}–plet of $SU(5)$ can belong to $\bf 6$}–plet, $\bf 15$}–plet and $\bf 20$}–plet of $SU(6)$. These $SU(6)$ representations have the following decomposition in terms of $SU(5)$ representations: ${\bf 6}={\bf 5}\oplus{\bf 1}$, ${\bf 15}={\bf 10}\oplus{\bf 5}$ and ${\bf 20}={\bf 10}\oplus{\bf\overline{10}}$.
Here we assume that the composite partners of the right-handed up- and down-type quarks ($U_{i}^{c}$ and $D_{i}^{c}$) stem from $\bf 20$}–plets and $\bf\overline{15}$}–plets of $SU(6)$ so that they belong to
$U^{c}_{i}\in\left({\bf\overline{3}},\,{\bf 3},\,-\dfrac{1}{3}\right)\,,\qquad D^{c}_{i}\in\Biggl{(}{\bf\overline{3}},\,{\bf\overline{3}},\,0\Biggr{)}\,.$ (8)
Then there are two types of the composite partners of the left–handed quarks ($Q_{1i}$ and $Q_{2i}$) which can be components of $\bf 15$}–plets and $\bf 20$}–plets. Therefore we have
$Q_{1i}\in\Biggl{(}{\bf 3},\,{\bf 3},\,0\Biggr{)}\,,\qquad Q_{2i}\in\left({\bf 3},\,{\bf\overline{3}},\,\dfrac{1}{3}\right)\,.$ (9)
Taking into account that the Higgs doublet is expected to be a part of the fundamental representation of $SU(6)$ (${\bf 6}_{H}$), so that $\Omega=\left({\bf 1},\,{\bf 3},\,\dfrac{1}{3}\right)$, the generalisation of the $SU(5)$ structure of the quark Yukawa interactions to the case of $SU(6)$ symmetry is given by [39]
$\mathcal{L}^{q}_{SU(6)}\sim Y^{u}_{ij}\,{\bf 20}(U^{c}_{i})\times{\bf 15}(Q_{1j})\times{\bf 6}_{H}+Y^{d}_{ij}\,{\bf 20}(Q_{2i})\times{\bf\overline{15}}(D^{c}_{j})\times{\bf\overline{6}}_{H}+h.c.\,.$ (10)
The $SU(6)$ structure of interactions (10) gives rise to the Yukawa couplings within the CHM3 that permit to induce the masses of all SM quarks at low energies.
It is also expected that the composite partners of the right-handed charged leptons originate from $\bf 15$}–plets of $SU(6)$ whereas two types of the partners of the left–handed leptons come from $\bf\overline{15}$}–plets and $\bf\overline{6}$}–plets. As a consequence one finds
$\begin{array}[]{c}E^{c}_{i}\in{\bf\overline{3}}^{i}_{+2/3}=\left({\bf 1},\,{\bf\overline{3}},\,\dfrac{2}{3}\right)\,,\qquad L_{1i}\in{\bf 3}^{i}_{-2/3}=\left({\bf 1},\,{\bf 3},\,-\dfrac{2}{3}\right)\,,\\ L_{2i}\in{\bf\overline{3}}^{i}_{-1/3}=\left({\bf 1},\,{\bf\overline{3}},\,-\dfrac{1}{3}\right)\,.\end{array}$ (11)
In order to ensure that the left–handed neutrino states gain masses which are much smaller than the charged lepton masses an approximate $Z_{2}$ discrete symmetry is imposed. All multiplets except ${\bf\overline{3}}^{i}_{-1/3}$ are even under $Z_{2}$ while ${\bf\overline{3}}^{i}_{-1/3}$ are $Z_{2}$ odd. Therefore the operators ${\bf\overline{3}}^{i}_{+2/3}\Omega^{\dagger}{\bf\overline{3}}^{i}_{-1/3}$ are strongly suppressed and only interactions
$\tilde{Y}^{e}_{ij}f({\bf\overline{3}}^{i}_{+2/3}\Omega)(\Omega^{\dagger}{\bf 3}^{j}_{-2/3})$ (12)
permit to reproduce in the strongly coupled sector the Yukawa couplings
$\mathcal{L}_{chl}\simeq Y^{e}_{ij}E^{c}_{i}(L_{1j}H^{c})+h.c.\,$, (13)
which give rise to non–zero masses of the charged leptons in the SM.
On the other hand the interactions $(L_{1i}H)(L_{1j}H)$, that could induce the masses of the left–handed neutrino, are forbidden by the $U(1)_{6}$ symmetry. The appropriate interactions in the strongly coupled sector
$\mathcal{L}_{nl}\simeq\dfrac{\varkappa_{ij}}{f}(L_{2i}H)(L_{2j}H)$ (14)
may come from the operators
$\tilde{\varkappa}_{ij}f({\bf\overline{3}}^{i}_{-1/3}\Omega)({\bf\overline{3}}^{j}_{-1/3}\Omega)$ (15)
which are not suppressed by neither $U(1)_{6}$ nor $Z_{2}$ symmetry. The $SU(6)$ structure, that leads to the interactions (12)-(15), can be written as
$\mathcal{L}^{e}_{SU(6)}\sim\eta_{ij}\Biggl{(}{\bf 15}(E^{c}_{i})\times{\bf\overline{6}}_{H}\Biggr{)}\times\Biggl{(}{\bf\overline{15}}(L_{1j})\times{\bf 6}_{H}\Biggr{)}+\xi_{ij}\Biggl{(}{\bf\overline{6}}(L_{2i})\times{\bf 6}_{H}\Biggr{)}\times\Biggl{(}{\bf\overline{6}}(L_{2j})\times{\bf 6}_{H}\Biggr{)}\,.$ (16)
The mixing between the elementary lepton doublets and their composite partners results in the SM lepton doublets $\ell_{i}$ as well as vectorlike fermion doublets $\tilde{L}_{1i}$ and $\tilde{L}_{2i}$ so that
[…]
where $c_{11i}\approx c_{22i}\approx 1$ while $s_{1i}$, $s_{2i}$, $s_{12i}$ and $s_{21i}$ are quite small. The observed mass hierarchy in the lepton sector implies that $s_{2i}\ll s_{1i}$. The smallness of $s_{2i}$ can be caused by the approximate discrete $Z_{2}$ symmetry. In this case $\tilde{L}_{2i}$ can be also much lighter than $f$. The scenarios with relatively light $\tilde{L}_{2i}$ are going to be considered in the next section.
The approximate $Z_{2}$ symmetry may also lead to the observed mass hierarchy in the lepton sector within the MCHM. Nevertheless from Eq. (5) it follows that the mass terms of the components of ${\bf 5}^{i}_{0}$ are not suppressed by the approximate $Z_{2}$ symmetry. As a consequence they can gain masses of the order of $f$.
3 Leptogenesis in the CHM3
Hereafter we explore the CHM3 with approximate $Z_{2}$ symmetry that leads to three generations of relatively light composite resonances $\tilde{L}_{2i}$. The SM singlet components $N_{i}$ of the corresponding $SU(3)$ multiplets ${\bf\overline{3}}^{i}_{-1/3}$ gain masses around $f$ through the interactions (15) while the $SU(2)_{W}$ doublet components of these multiplets are substantially lighter than $f$. This scenario is realized in Nature if the compositeness scale $f\gtrsim 10\,\mbox{TeV}$. In this case the masses of $\tilde{L}_{2i}$ may vary between $1-2\,\mbox{TeV}$ so that it might be possible to discover such fermion states in the near future at the LHC. The couplings of these composite resonances to the SM leptons are suppressed because they are determined by the small parameters $s_{2i}$.
Now suppose that the weakly–coupled sector includes an additional elementary Majorana fermion $n$ which mixes with the SM singlet resonances $N_{i}$. The corresponding mixing parameters are expected to be of order of $s_{2i}\lesssim 10^{-5}$. If $n$ and $N_{i}$ have masses around $10\,\mbox{TeV}$ they can decay into leptons and Higgs doublet inducing lepton asymmetry. We further assume that $N_{i}$ are somewhat heavier than the elementary fermion $n$. The part of the CHM3 Lagrangian, which describes the interactions of $n=N_{0}$ and $N_{i}$ with the SM leptons, $\tilde{L}_{2i}$ and Higgs doublet $H$, can be presented as a sum
$\mathcal{L}_{N}=g_{ix}\ell_{i}HN_{x}+h_{jx}\tilde{L}_{2j}HN_{x}+h.c.\,$, (18)
where $i,j=1,2,3$ and $x=0,1,2,3$. Since $\ell_{i}$ and $N_{0}$ are mainly elementary fields $|g_{i0}|\ll|g_{ij}|$, $|g_{ij}|\sim|h_{j0}|\ll|h_{ij}|$. At the same time the Yukawa couplings $h_{ij}$ are not suppressed by the $Z_{2}$ symmetry and therefore $|h_{ij}|\gtrsim 0.1$.
To simplify our analysis, we ignore the couplings $g_{ix}$, neglect the masses of $\tilde{L}_{2i}$ and set $h_{20}=h_{30}=0$. Then the process of the lepton asymmetry generation is controlled by only one CP (decay) asymmetry
$\varepsilon_{0}\simeq\dfrac{\Gamma_{L_{21}}-\Gamma_{\overline{L}_{21}}}{\left(\Gamma_{L_{21}}+\Gamma_{\overline{L}_{21}}\right)}$ (19)
which appears on the right–hand side of Boltzmann equations. Here $\Gamma_{L_{21}}$ and $\Gamma_{\overline{L}_{21}}$ are partial decay widths of $n\to\tilde{L}_{21}+H$ and $n\to\overline{\tilde{L}}_{21}+H^{*}$. At the tree level the decay asymmetry (19) vanishes because
$\Gamma_{L_{21}}=\Gamma_{\overline{L}_{21}}=\dfrac{|h_{10}|^{2}}{16\pi}\,M_{n}\,.$ (20)
In Eq. (20) $M_{n}$ is the mass of the Majorana fermion $n$. The CP violation gives rise to the non–zero value of $\varepsilon_{0}$ that appears because of the interference between the tree–level amplitude of the decays of $n$ and one–loop corrections to it. For $h_{1x}=|h_{1x}|e^{i\varphi_{x}}$ and real values of $M_{n}$ and $M_{j}$ the calculation of one–loop diagrams gives [42]
$\varepsilon_{0}\simeq\dfrac{1}{8\pi}\Biggl{[}\sum_{j=1,2,3}|h_{1j}|^{2}g\left(\frac{M^{2}_{j}}{M_{n}^{2}}\right)\sin 2\Delta\varphi_{j}\Biggr{]}\,,\qquad\qquad\Delta\varphi_{j}=\varphi_{j}-\varphi_{0}\,$, (21)
$g(x)=\sqrt{x}\biggl{[}\dfrac{1}{1-x}+1-(1+x)\ln\dfrac{1+x}{x}\biggr{]}\,$,
where $M_{j}$ are the masses of the composite resonances $N_{j}$. When CP invariance is preserved, i.e. $\varphi_{x}\to 0$, the decay asymmetry (21) vanishes. The absolute value of $\varepsilon_{0}$ reaches its maximum for $\Delta\varphi_{j}=\Delta\varphi=\pm\pi/4$.
It is convenient to introduce $Y_{\Delta B}$ which is the baryon asymmetry relative to the entropy density. The observed value of $Y_{\Delta B}$ is given by
$Y_{\Delta B}=\dfrac{n_{B}-n_{\bar{B}}}{s}\biggl{|}_{0}=(8.75\pm 0.23)\times 10^{-11}\,.$ (22)
Using an approximate formula [42]
$Y_{\Delta B}\sim 10^{-3}\,\varepsilon_{0}\,\eta_{0}\,$, (23)
one can estimate the induced baryon asymmetry. In Eq. (23) $\eta_{0}$ is an efficiency factor that varies from 0 to 1. Without washout effects the decays of the Majorana fermion $n$ would result in $\eta_{0}=1$. The washout processes reduce the generated baryon asymmetry by the factor $\eta_{0}$. Here we concentrate on the so–called strong washout scenario when the efficiency factor $\eta_{0}$ can be estimated as
$\eta_{0}\simeq H(T=M_{1})/\Gamma\,$$\qquad\qquad\qquad$
$\Gamma=\Gamma_{L_{21}}+\Gamma_{\overline{L}_{21}}\,,\qquad\qquad H=1.66\,g_{*}^{1/2}\dfrac{T^{2}}{M_{Pl}}\,$, (24)
where $H$ is the Hubble expansion rate and $g_{*}=n_{b}+\dfrac{7}{8}\,n_{f}$ is the number of relativistic degrees of freedom in the thermal bath. In the scenario under consideration $g_{*}=128.75$.
From Eqs. (24) it follows that $\eta_{0}$ diminishes with increasing $|h_{10}|$. The corresponding dependence is shown in Fig. 1a. In Eq. (23) both the decay asymmetry $\varepsilon_{0}$ and the efficiency factor are smaller than unity. Thus the appropriate baryon asymmetry can be generated only if $\eta_{0}\gtrsim 10^{-7}$. The results of the numerical analysis presented in Fig. 1a demonstrate that for $M_{n}\simeq 10\,\mbox{TeV}$ this condition can be fulfilled when $|h_{10}|\lesssim 0.001$. Otherwise the absolute value of $Y_{\Delta B}$ tends to be too small. Within the CHM3 the value of $|h_{10}|$ is expected to be of the order of $s_{2i}$ because fermion $n$ is mainly elementary state, i.e. $|h_{10}|\lesssim 10^{-5}$.
(a)                        (b)
Figure 1: Logarithm (base 10) of the absolute values of the efficiency factor $\eta_{0}$ and decay asymmetry $\varepsilon_{0}$ for $h_{20}=h_{30}=0$. In (a) the absolute value of $\eta_{0}$ is given as a function of logarithm (base 10) of $|h_{10}|$ for $M_{n}=5\,\mbox{TeV}$ (dashed–dotted line), $M_{n}=10\,\mbox{TeV}$ (solid line) and $M_{n}=15\,\mbox{TeV}$ (dashed line). In (b) the absolute value of $\varepsilon_{0}$ is presented as a function of logarithm (base 10) of $\Delta\varphi_{1}=\Delta\varphi_{2}=\Delta\varphi_{3}=\Delta\varphi$ for $M_{n}=10\,\mbox{TeV}$, $M_{1}=12\,\mbox{TeV}$, $M_{2}=14\,\mbox{TeV}$, $M_{3}=16\,\mbox{TeV}$, $|h_{11}|=|h_{12}|=|h_{13}|=|h|$, $|h|=0.1$ (solid line) and $|h|=0.3$ (dashed line).
Since in the scenario under consideration the efficiency factor is not negligibly small, i.e. $\eta_{0}\sim 0.01$, the absolute value of $Y_{\Delta B}$ is determined by the decay asymmetry $\varepsilon_{0}$. From Eq. (21) one can see that $\varepsilon_{0}$ is set by $|h_{1j}|$ and CP violating phases $\Delta\varphi_{j}$. It also depends on the ratio of masses of Majorana fermions $M_{j}/M_{n}$. Because the Yukawa couplings of $N_{j}$ to $L_{2i}$ and the Higgs doublet are not suppressed by the approximate $Z_{2}$ symmetry, all $|h_{ij}|$ should be relatively large, i.e. $|h_{ij}|\gtrsim 0.1$. Here we fix $M_{n}=10\,\mbox{TeV}$, $M_{1}=12\,\mbox{TeV}$, $M_{2}=14\,\mbox{TeV}$, $M_{3}=16\,\mbox{TeV}$, $|h_{11}|=|h_{12}|=|h_{13}|=h$ and $\Delta\varphi_{1}=\Delta\varphi_{2}=\Delta\varphi_{3}=\Delta\varphi$. In Fig. 1b the dependence of $|\varepsilon_{0}|$ on $\Delta\varphi$ for different values of $h$ is shown. The absolute value of the decay asymmetry grows monotonically with increasing of $|h|$. The results presented in Fig. 1b indicate that for $h\gtrsim 0.1$ and $\eta_{0}\sim 0.01$ the observed baryon asymmetry can be obtained for small CP violating phases, i.e. $\Delta\varphi\ll 0.01$.
4 Conclusions
In this paper we have discussed the 331 composite Higgs model (CHM3) in which the strongly interacting sector possesses approximate $SU(3)_{C}\times SU(3)\times U(1)_{6}$ symmetry. This global symmetry may originate from $SU(6)$ subgroup of $E_{6}$. It is expected that near the scale $f\sim 10\,\mbox{TeV}$ the approximate $SU(3)\times U(1)_{6}$ symmetry is broken down to the $SU(2)_{W}\times U(1)_{Y}$ subgroup giving rise to five pNGBs. These pNGB states form Higgs doublet and one SM singlet scalar. The composite partners of the SM quarks and SM leptons belong to fundamental and antifundamental representations of $SU(3)$ with different $U(1)_{6}$ charges.
We argued that the observed mass hierarchy in the lepton sector can be caused by the approximate discrete $Z_{2}$ symmetry. Such $Z_{2}$ symmetry may also lead to three generations of relatively light composite neutral fermions and fermions with charges $\pm 1$. These resonances compose $SU(2)_{W}$ doublets $\tilde{L}_{2i}$ and can have masses within the $1-2\,\mbox{TeV}$ range so that they can be discovered at the LHC in the near future. The couplings of $\tilde{L}_{2i}$ to the SM leptons are quite suppressed because they are defined by the small parameters $s_{2i}\lesssim 10^{-5}$. It was shown that in this case the observed baryon asymmetry can be induced if the particle spectrum of the CHM3 contains an additional elementary Majorana fermion $n$ with mass around $10\,\mbox{TeV}$. The corresponding lepton asymmetry is generated due to the out–of equilibrium decays of $n$ into $\tilde{L}_{2i}$ and Higgs doublet even when CP is almost preserved.
Acknowledgements
RN is grateful to M. G. Belyakova, E. E. Boos, S. V. Demidov, D. S. Gorbunov, D. I. Kazakov, N. V. Krasnikov, D. G. Levkov and M.I. Vysotsky for fruitful discussions.